# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ VIII Outlook
We further derived the response of any (generally path-wise) observable to a general (incl. thermal) perturbation. Finally, we established the continuum limit to achieve the complete unification of diffusion and jump dynamics. Our results ultimately place functionals of diffusion and jump processes on an equal footing on the level of individual stochastic realizations, and hence achieve a “contraction” of two until now disjoint frameworks in time-average statistical mechanics.
@@ Acknowledgments
The financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to AG) and the German Research Foundation (DFG) through the Heisenberg Program (grant GO 2762/4-1 to AG) is gratefully acknowledged.
@@ A.1 Proof of Noise-Time Correlation Lemma
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=$ $\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
$+\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle,$ (164)
@@
$\mathds{1}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
@@
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=\mathds{1}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle.$ (166)
@@
$\mathrm{d}\varepsilon_{xy}(\tau)=\mathds{1}_{x}(\tau)\begin{cases}1-r_{xy}\mathrm{d}\tau&\text{transition},\\ -r_{xy}\mathrm{d}\tau&\text{no transition}.\end{cases}$ (167)
@@
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$ $=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\sum_{\mathrm{d}{\varepsilon}_{xy}(\tau)}p(\mathrm{d}{\varepsilon}_{xy}(\tau))\mathrm{d}{\varepsilon}_{xy}(\tau)P(i,\tau^{\prime}|\gamma(\mathrm{d}{\varepsilon}_{xy}(\tau)),\tau+\mathrm{d}\tau)p_{x}(\tau)$
$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}p_{x}(\tau)r_{xy}\mathrm{d}\tau\left(1-r_{xy}\mathrm{d}\tau\right)\left[P(i,\tau^{\prime}|y,\tau+\mathrm{d}\tau)-P(i,\tau^{\prime}|x,\tau+\mathrm{d}\tau)\right]$
$=\mathds{1}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}\mathrm{d}\tau\,.$ (169)
@@ A.2.1 Jump-Time Correlations
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle}{r_{xy}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (171)
@@ A.2.2 Jump-Jump Correlations
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle}{r_{xy}r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ $\delta_{ix}\delta_{jy}\delta(\tau-\tau^{\prime})p_{x}(\tau)$
$+\mathds{1}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+\mathds{1}_{\tau\geq\tau^{\prime}}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime}).$ (173)
@@ A.3.1 Continuous Space: Langevin to Fokker-Planck
To highlight the analogy between continuous- and discrete space dynamics, we first recapitulate how the Fokker-Planck equation follows directly from the Langevin Equation in Eq. (31). The calculation is well ḱnown (see e.g., Ref. [132]). Let $K(x)$ be an arbitrary test function that obeys the boundary conditions. Considering the expectation $\langle K(\bm{x}(t))\rangle=\int\mathrm{d}\bm{x}K(\bm{x})P(\bm{x},t)$ for the dummy function at some time $t$, the temporal derivative can be evaluated using Itô’s lemma [132]
@@
$=\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t).$
@@
$=$ $\int\mathrm{d}\bm{x}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]P(\bm{x},t)$ (176)
@@ A.3.2 Discrete Space: Jumps to Master Equation
To show how the master equation emerges from the stochastic equations of motion in Eq. (38), we first need to consider the indicator function $\mathds{1}_{x}(\tau)$ of being in state $x$ at time $\tau$. Consider $\mathds{1}_{x}(\tau+\mathrm{d}\tau)=\mathds{1}_{x}(\tau)+\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$, where the derivative should be understood as
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)=\lim_{\mathrm{d}\tau\to 0^{+}}\frac{\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)}{\mathrm{d}\tau}.$ (178)
By the non-explosive property of MJP [169] a finite number of jumps occur in any finite time interval with probability 1. Therefore, we do not need to consider any higher order correction, as the probability of a second transition occurring in an infinitesimal increment vanishes. The term $\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau$ can therefore take on values
@@
$\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau=\sum_{y\neq x}\left(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau)\right),$ (180)
because at most one jump differential will be non-zero as $\mathrm{d}\tau\to 0$. The probability of being in a state $x$ at time $\tau$ is $p_{x}(\tau)=\langle\mathds{1}_{x}(\tau)\rangle$. Therefore,
$p_{x}(\tau+\mathrm{d}\tau)-p_{x}(\tau)$ $=\left\langle\mathds{1}_{x}(\tau+\mathrm{d}\tau)-\mathds{1}_{x}(\tau)\right\rangle$
$=\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau}\mathds{1}_{x}(\tau)\mathrm{d}\tau\right\rangle$ (181)
$=\left\langle\sum_{y\neq x}(\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau))\right\rangle$
$=\mathrm{d}\tau\sum_{y\neq x}\left(r_{yx}p_{y}(\tau)-r_{xy}p_{x}(\tau)\right)$
@@
$=$ $\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$
@@
$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}\mathds{1}_{\tau<\tau^{\prime}}$ (182)
@@
$-\left.\int_{0}^{t}\mathrm{d}\tau\sum_{x}p_{x}(\tau)\partial_{\tau}\left\{P(i,\tau^{\prime}|x,\tau)\mathds{1}_{\tau<\tau^{\prime}}\right\}\right]$
@@
where we use that $\partial_{\tau}P(i,\tau^{\prime}|x,\tau)=-\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$, $\partial_{\tau}\mathds{1}_{\tau<\tau^{\prime}}=-\partial_{\tau^{\prime}}\mathds{1}_{\tau<\tau^{\prime}}$, and $\int_{0}^{\tau^{\prime}}\mathrm{d}\tau\mathds{1}_{\tau<\tau^{\prime}}=\tau^{\prime}$. With another integration by parts, the correlator finally becomes
@@ A.6 Long-time limit of driven ring
Here we present details on the large-$t$ limit of the stationary correlation bound in Eq. (123) applied to the driven ring Sec. IV.1 with the assumption that $\epsilon>1$. We start by realizing that $\Delta\bm{V}\equiv\bm{V}^{\mathrm{A}}-\langle{V}^{\mathrm{A}}\rangle\bm{1}=\frac{1}{2}(1,1,-1,-1)^{T}$. Hence, we can write
@@ A.7 Physically Motivated Rates - Pseudo Potential Form
$\partial_{\delta}O_{t}=$ $\lim_{\delta\to 0}\frac{1}{\delta}\left(\left\langle O(t)\frac{\mathrm{d}\mathds{P^{\delta}}}{\mathrm{d}\mathds{P}}\right\rangle-\langle O(t)\rangle\right)$
@@ A.8 Perturbations of Path Observables
$O(t)=\int_{s=0}^{s=t}\bm{g}^{T}\mathrm{d}\bm{\tau}(s),$ (208)
with some state function vector $\bm{g}$. The response is
@@
$\partial_{\delta}O_{t}=\sum_{i,x,y\neq x}\int_{0}^{t}\mathrm{d}\tau$ $\int_{0}^{t}\mathrm{d}s\mathds{1}_{s<\tau}g_{i}\tilde{r}_{xy}p_{x}(s)$ (214)
@@
Lastly, although not explicitly written, the matrix $\bm{b}$ and vector $\bm{g}$ may also depend on time. Equations (205) and (213) are written to readily accommodate for this.

### v2

@@ IX Outlook
We further derived the response of any (generally path-wise) observable to a general (incl. thermal) perturbation and constructed a corresponding response-function formalism for single-time as well as path-wise observables. Subsequently, we established the continuum limit to achieve the complete unification of diffusion and jump dynamics. Lastly, we derived the classical analogue of the Belavkin equation and commented on the analogy with open quantum systems. Our results ultimately place functionals of diffusion and jump processes on an equal footing on the level of individual stochastic realizations, and hence achieve a “contraction” of two until now disjoint frameworks in time-average statistical mechanics. Our newly developed framework has recently already been applied in [174].
@@ Acknowledgments
We thank Jorge ”Jefe” Tabanera-Bravo for discussions on quantum systems. The financial support from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program (Grant Agreement No. 101086182 to AG) and the German Research Foundation (DFG) through the Heisenberg Program (grant GO 2762/4-1 to AG) is gratefully acknowledged.
@@ A.1 Proof of Noise-Time Correlation Lemma
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle=$ ${\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle,$ (196)
@@
${\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$
@@
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle.$ (198)
@@
$\mathrm{d}\varepsilon_{xy}(\tau)={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}(\tau)\begin{cases}1-r_{xy}\mathrm{d}\tau&\text{transition},\\ -r_{xy}\mathrm{d}\tau&\text{no transition}.\end{cases}$ (199)
@@
$\langle\mathrm{d}{\varepsilon}_{xy}(\tau)\mathrm{d}{\tau}_{i}(\tau^{\prime})\rangle$ $={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\sum_{\mathrm{d}{\varepsilon}_{xy}(\tau)}p(\mathrm{d}{\varepsilon}_{xy}(\tau))\mathrm{d}{\varepsilon}_{xy}(\tau)P(i,\tau^{\prime}|\gamma(\mathrm{d}{\varepsilon}_{xy}(\tau)),\tau+\mathrm{d}\tau)p_{x}(\tau)$
$={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}p_{x}(\tau)r_{xy}\mathrm{d}\tau\left(1-r_{xy}\mathrm{d}\tau\right)\left[P(i,\tau^{\prime}|y,\tau+\mathrm{d}\tau)-P(i,\tau^{\prime}|x,\tau+\mathrm{d}\tau)\right]$
$={\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\mathrm{d}\tau^{\prime}\left[P(i,\tau^{\prime}|y,\tau)-P(i,\tau^{\prime}|x,\tau)\right]p_{x}(\tau)r_{xy}\mathrm{d}\tau\,.$ (201)
@@ A.2.1 Jump-Time Correlations
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}\tau_{i}(\tau^{\prime})\rangle}{r_{xy}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ ${\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}P(x,\tau|i,\tau^{\prime})p_{i}(\tau^{\prime}).$ (203)
@@ A.2.2 Jump-Jump Correlations
$\frac{\langle\mathrm{d}n_{xy}(\tau)\mathrm{d}n_{ij}(\tau^{\prime})\rangle}{r_{xy}r_{ij}\mathrm{d}\tau\mathrm{d}\tau^{\prime}}=$ ${\color[rgb]{0,0,0}r_{ij}}\delta_{ix}\delta_{jy}\delta(\tau-\tau^{\prime})p_{x}(\tau)$
$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}P(i,\tau^{\prime}|y,\tau)p_{x}(\tau)$
$+{\color[rgb]{0,0,0}\mathbb{1}}_{\tau\geq\tau^{\prime}}P(x,\tau|j,\tau^{\prime})p_{i}(\tau^{\prime}).$ (205)
@@ A.3.1 Continuous Space: Langevin to Fokker-Planck
To highlight the analogy between continuous- and discrete space dynamics, we first recapitulate how the Fokker-Planck equation follows directly from the Langevin Equation in Eq. (39). The calculation is well known (see e.g., Ref. [51]). Let $K(x)$ be an arbitrary test function that is sufficiently smooth and obeys the boundary conditions. Considering the expectation $\langle K(\bm{x}(t))\rangle=\int\mathrm{d}\bm{x}K(\bm{x})P(\bm{x},t)$ for the dummy function at some time $t$, the temporal derivative can be evaluated using Itô’s lemma [51]
@@
$=\int\mathrm{d}\bm{x}{\color[rgb]{0,0,0}P(\bm{x},t)}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right].$
@@
$=$ $\int\mathrm{d}\bm{x}{\color[rgb]{0,0,0}P(\bm{x},t)}\left[\bm{F}(\bm{x})\cdot\nabla K(\bm{x})+\nabla\cdot\bm{D}\nabla K(\bm{x})\right]$ (208)
@@ A.3.2 Discrete Space: Jumps to Master Equation
To show how the master equation emerges from the stochastic equations of motion in Eq. (46), we first need to consider the indicator function $\color[rgb]{0,0,0}\delta_{x_{\tau}x}$} of being in state $x$ at time $\tau$. Consider ${\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}={\color[rgb]{0,0,0}\delta_{x_{\tau}x}}+\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau+\mathcal{O}(\mathrm{d}\tau^{2})$, where the derivative should be understood as
$\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}=\lim_{\mathrm{d}\tau\to 0^{+}}\frac{{\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}-{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}}{\mathrm{d}\tau}.$ (210)
By the non-explosive property of MJP [21] a finite number of jumps occur in any finite time interval with probability 1. Therefore, we do not need to consider any higher order correction, as the probability of a second transition occurring in an infinitesimal increment vanishes. The term $\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau$ can therefore take on values
@@
$\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau=\sum_{y\neq x}{\color[rgb]{0,0,0}[}\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau){\color[rgb]{0,0,0}]},$ (212)
because at most one jump differential will be non-zero as $\mathrm{d}\tau\to 0$. The probability of being in a state $x$ at time $\tau$ is $p_{x}(\tau)=\langle{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\rangle$. Therefore,
$p_{x}(\tau+\mathrm{d}\tau)-p_{x}(\tau)$ $=\left\langle{\color[rgb]{0,0,0}\delta_{x_{\tau+\mathrm{d}\tau}x}}-{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\right\rangle$
$=\left\langle\frac{\mathrm{d}}{\mathrm{d}\tau}{\color[rgb]{0,0,0}\delta_{x_{\tau}x}}\mathrm{d}\tau\right\rangle$ (213)
$=\left\langle\sum_{y\neq x}{\color[rgb]{0,0,0}[}\mathrm{d}n_{yx}(\tau)-\mathrm{d}n_{xy}(\tau){\color[rgb]{0,0,0}]}\right\rangle$
$=\mathrm{d}\tau\sum_{y\neq x}{\color[rgb]{0,0,0}[}r_{yx}p_{y}(\tau)-r_{xy}p_{x}(\tau){\color[rgb]{0,0,0}]}$
@@
$=$ $\frac{1}{2}\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$
@@
$=$ $-\int_{0}^{t}\mathrm{d}\tau^{\prime}\int_{0}^{t}\mathrm{d}\tau\sum_{i,j}\kappa_{ij}(\tau^{\prime})r_{ij}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$ (214)
@@
$-\left.\int_{0}^{t}\mathrm{d}\tau\sum_{x}p_{x}(\tau)\partial_{\tau}\left\{P(i,\tau^{\prime}|x,\tau){\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}\right\}\right]$
@@
where we use that $\partial_{\tau}P(i,\tau^{\prime}|x,\tau)=-\partial_{\tau^{\prime}}P(i,\tau^{\prime}|x,\tau)$, $\partial_{\tau}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=-\partial_{\tau^{\prime}}{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}$, and $\int_{0}^{\tau^{\prime}}\mathrm{d}\tau{\color[rgb]{0,0,0}\mathbb{1}}_{\tau<\tau^{\prime}}=\tau^{\prime}$. With another integration by parts, the correlator finally becomes
@@ A.6 Long-time limit of driven ring
Here we present details on the large-$t$ limit of the stationary correlation bound in Eq. (131) applied to the driven ring Sec. IV.1 with the assumption that $\epsilon>1$. We start by realizing that $V^{\Delta}_{\tau}=V_{\tau}-\langle V_{\tau}\rangle=\bm{\delta}(\tau)^{T}\Delta\bm{V}$ with $\bm{\delta}(\tau)^{T}=(\delta_{x_{\tau}1},\delta_{x_{\tau}2},\delta_{x_{\tau}3},\delta_{x_{\tau}4})$ and $\Delta\bm{V}\equiv\bm{V}^{\mathrm{A}}-\langle{V}^{\mathrm{A}}\rangle\bm{1}=\frac{1}{2}(1,1,-1,-1)^{T}$. Hence, we can write
@@ A.7 Physically Motivated Rates - Pseudo Potential Form
$\partial_{\delta}O_{t}=$ $\lim_{\delta\to 0}\frac{1}{\delta}\left(\left\langle O(t)\frac{\mathrm{d}\mathbb{P^{\delta}}}{\mathrm{d}{\color[rgb]{0,0,0}\mathbb{P}}}\right\rangle-\langle O(t)\rangle\right)$
@@ A.8 Perturbations of Path Observables
$O(t)=\int_{s=0}^{s=t}{\color[rgb]{0,0,0}g_{s}\mathrm{d}s},$ (240)
with some state function $g_{\tau}\equiv\sum_{k}\delta_{x_{\tau}k}g_{k}$. The response is
@@
$\partial_{\delta}O_{t}=\sum_{i,x,y\neq x}\int_{0}^{t}\mathrm{d}\tau$ $\int_{0}^{t}\mathrm{d}s{\color[rgb]{0,0,0}\mathbb{1}}_{s<\tau}g_{i}\tilde{r}_{xy}p_{x}(s)$ (246)
@@
Lastly, although not explicitly written, the matrix $\bm{b}$ and functions $g_{k}$ may also depend on time. Equations (237) and (245) are written to readily accommodate for this.